samuel is 12 years older than his brother james. the product of their ages is 133. how old is each brother?
Answer by TimothyLamb(4379)

x = sam age now
y = jim age now
---
x = y + 12
xy = 133
---
xy = 133
(y + 12)y = 133
yy + 12y - 133 = 0
---
the above quadratic equation is in standard form, with a=1, b=12 and c=-133
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 12 -133
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
y = 7
y = -19
---
the negative root doesn't fit the problem statement, so use the positive root:
---
y = 7
x = y + 12
x = 7 + 12
x = 19
---
answer:
y = jim age now = 7
x = sam age now = 19
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php